OpenStudy (anonymous):

Rearrange the equation x=F(x) into the form f(x)=0, where f is a polynomial function. (I'm writing it up, so hold on)

OpenStudy (anonymous):

\[X_{0}=0, ~~~X_{r+1}= \sqrt[11]{x_{r}~^7-6}\]

OpenStudy (anonymous):

raise both sides to the power of 11 subtract x^7 add 6

OpenStudy (anonymous):

\[x=\sqrt[11]{x^7-6}\] \[x^{11}=x^7-6\] \[x^{11}-x^7+6=0\]

OpenStudy (anonymous):

Thanks so much :)

OpenStudy (anonymous):

easy once you see it right?

OpenStudy (anonymous):

Yes! The next step is to use the iteration, with the given initial approximation x_0 to find the terms of the sequence x_0, x_1,..... As far as x_5?

OpenStudy (anonymous):

\[x_0, x_1,..... As ~far~ as~~~ x_5\]?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

i think for this you use \[x_{n+1}=F(x_n)\]

OpenStudy (anonymous):

Ok, and how do you apply the formula?

OpenStudy (anonymous):

need a calculator for sure

OpenStudy (anonymous):

you have \[x_{n+1}=\sqrt[11]{x^7-6},x_0=0\] so \[x_1=\sqrt[11]{0^7-6}=\sqrt[11]{-6}\] whatever that is

OpenStudy (anonymous):

i get \[x_1=-1.1769\] rounded

OpenStudy (anonymous):

It shows the answer in decimals. Yes, that's the answer. but I can't get that on my calculator? How did you get it?

OpenStudy (anonymous):

and then \[x_2=\sqrt[11]{(-1.1769)^7-6}\]etc

OpenStudy (anonymous):

depends on your calculator

OpenStudy (anonymous):

Casio scientific

OpenStudy (anonymous):

i used a cheap one, but if you have a nice one you can take \[(-6)^{\frac{1}{11}}\]

OpenStudy (anonymous):

Mine gives -1.0859? Why?

OpenStudy (anonymous):

maybe i am wrong

OpenStudy (anonymous):

no i think i am right

OpenStudy (anonymous):

Yes. I got it now. Thanks :) I made a mistake... I shouldn't have added the - until after I calculated the square.

OpenStudy (anonymous):

ok good! you have four more to do good luck

OpenStudy (anonymous):

Thanks again! :D